# BA-SAM: Scalable Bias-Mode Attention Mask for Segment Anything Model

### 04 Jan 2024 v1

@@ 4.2 Bias-Mode Attention Mask
Another challenge is that changes in token sequence length will lead to variations in positional encoding. It is important to ensure the insensitivity of the model when such positional encoding variations occur during testing.
One possible way is absolute encoding without trainable parameters, such as Sinusoidal [62]. It requires the position encoding to have strong local-to-global inference capabilities. As mentioned by Su [60], this resembles Taylor series approximations: by knowing the values of several derivatives at a particular point, effective estimates can be made within a neighborhood. Nevertheless, this assumes that the given function has high-order smoothness (higher-order derivatives exist and are bounded). Commonly used positional encodings are often combined with trigonometric functions, essentially high-frequency oscillatory functions with periodicity. However, these methods fail to satisfy the requirement of bounded high-order derivatives, making it less accurate to estimate the extrapolated results.
Another potential approach is using local attention, which constrains the model’s field of view and remains insensitive to variations in token sequence length. However, local attention is typically implemented using a local window, necessitating modifications to the SAM structure, which requires re-training from scratch and is unacceptable for deployment in the downstream tasks.
To this end, we propose to enable the attention layer to put more focus on neighboring tokens of the current token. As such, even with an increase in the length of a token sequence, each token is scarcely affected by the untrained tokens from distant positions. In particular, we design a simple yet effective bias-mode mask. We introduce a bias after the query-key dot product.
As shown in Figure 3, this mask exhibits a bias specified on the distance between the query-key pairs (i.e., $q\cdot k$). We expect that this proposed mask imposes penalties on attention scores between distant query-key pairs, and the penalty increases as the distance between a key $q$ and a query $k$ grows. To this end, we simply define the bias as the form of $b_{i,j}=\beta|i-j|$.
@@
We further discuss the setting of $\beta$ based on different cases. When conducting zero-shot transfer learning without fine-tuning, we set $\beta$ to a static, non-learned fixed value. The specific value setting will be discussed in the experimental section (Sec. 5). When fine-tuning is required, we make $\beta$ trainable. Since our Bias-Mode Attention Mask is lightweight relative to the model structure, it incurs negligible training time overhead.
@@ 4.3 BA-SAM Model
As shown in Figure 3, our BA-SAM is simple to implement, and can be seamlessly integrated into SAM [38] and its variants. Specifically, our design involves a new scaling factor (NSF) for the attention layer and a bias-mode attention mask (BM-AM). Our method does not involve any alterations to the model structure and is suited to both fine-tuning and non-fine-tuning cases. As for the fine-tuning case, it introduces negligible computational overhead as BM-AM incurs a very small amount of computation.
@@ 5 Experiments
Table 1: Performance comparisons in varying image resolutions. We employed the widely-used MAE (Mean Absolute Error) score. Lower MAE scores indicate better model accuracy. $\Delta$diff denotes the performance degradation due to resolution changes. Compared to the SAM [38] and MobileSAM [76] baselines, our BA-SAM achieves smaller degradation when encountering token sequences length changes. The best MAE performance is highlighted in bold, and the smallest performance degradation is underlined.
We first describe the datasets and implementation details in Sec. 5.1. Then, we compare BA-SAM’s performance with state-of-the-art methods across multiple tasks in Sec. 5.2. Finally, extensive ablation studies and analysis are conducted in Sec. 5.3. to demonstrate the effectiveness.
@@ 5.1 Datasets and Implementations
Datasets. For a comprehensive evaluation of BA-SAM, we conduct extensive experiments on a wide range of segmentation tasks, i.e., salient object segmentation [18, 15, 20], complex object segmentation [19], skin lesion segmentation [93], camouflaged object detection [17], which correspond to four datasets: DUTS [64], DIS-TE4 [53], ISIC [12] and COD10K [16]. Besides, we also verify it on the challenging COCO [45] instance segmentation benchmark. More details about datasets can be referred to in the supplementary material.
Implementation Details. In zero-shot settings that do not require fine-tuning, we use the original SAM [38] backbone. For fine-tuning scenarios, we employ MobileSAM [76] as the backbone. MobileSAM [76] is a SAM variant with a structure similar to ViT-Tiny [72], and further details can be found in [76]. MobileSAM [76] uses a ViT-H-based SAM as a teacher network for distillation, ultimately achieving competitive accuracy compared to the original SAM but with significantly fewer parameters. For various object segmentation tasks, a random point is extracted from the ground truth as the prompt input during the fine-tuning phase. For instance segmentation, we use the ViT-B [14] backbone and the state-of-the-art detector Deformable-DETR  [91] trained on the COCO [45] dataset with Swin-L [47] backbone as box prompt generator. More details are provided in the supplementary material. The code will be released at this link.
@@ 5.2 Results
Results of Various Object Segmentation Tasks: Table 1 demonstrates the effectiveness of our approach across four diverse segmentation datasets. $\Delta$diff denotes the value of the performance degradation due to resolution changes during the inference. The upper and lower parts of the table indicate the results without and with fine-tuning, respectively due to input resolution variations. The best MAE performance is highlighted in bold, and the smallest degradation is underlined. We have three observations: Firstly, our proposed BA-SAM consistently outperforms both SAM [38] and MobileSAM [76] baselines on all four datasets. This is mainly because these baselines do not consider the issue of varying image resolutions. In contrast, our presented scaling factor and bias-mode attention mask explicitly handle this issue and further alleviate the performance degradation. Secondly, when testing on higher resolutions than the training size, SAM [38] and MobileSAM [76] baselines show less desirable results than the original image size. In contrast, our BA-SAM incurs significantly less performance drop in different datasets. Thirdly, during the experiment, we observe negligible computational overhead, whether fine-tuning is applied or not, which supports the claim in the method section. See Sec. 5.3 for details.
Results of Instance Segmentation: In Table 3, we evaluate the performance of our method on the COCO [45] instance segmentation benchmark. For a fair comparison, all experiments are conducted in a zero-shot manner, with the same initialization parameters for the comparative methods and without the use of any additional training data. Our BA-SAM consistently outperforms SAM [38] and MobileSAM [76] baselines, demonstrating better zero-shot generalization capability on instance segmentation.
Comparisons with State-of-the-Art Methods: To further demonstrate the superiority of the effectiveness and generalizability of our method, we compare the state-of-the-art approaches in Table 2. From the table, we have two following observations: Firstly, all the state-of-the-art approaches [57, 77, 46, 92, 81, 53, 37, 16] show less-desirable performances in each dataset. Instead, our BA-SAM (specialized models) consistently outperforms these methods when fine-tuned on each downstream dataset. Secondly, almost all of these state-of-the-art techniques are specifically designed for one task and cannot be generalized well to other tasks. Due to the strong zero-shot generalization capability of SAM [38], our proposed BA-SAM can also be employed as a generalized model, which fine-tunes with all of these downstream datasets in a unified and shared model. Importantly, unlike [46, 57, 77, 12], we eliminate the need for employing additional techniques to further enhance the performance. As shown in Table 2, our generalized model also consistently promotes the performance of SAM on all datasets, demonstrating its remarkable generalizability.
@@ 5.3 Ablation Study and Analysis
Ablations Studies of Each Component. Table 4 summarizes the effect of each designed component on the settings with and without fine-tuning, respectively. The baseline means using the MobileSAM [76] as the base network that uses the vanilla scaling factor (VSF) in the attention layer [62]. New Scaling Factor and Bias-Mode Attention Mask are abbreviated as NSF and BM-AM, respectively. From the table, we observe that NSF could achieve superior performances than the baseline with VSF. This is because the vanilla attention in SAM [38] and MobileSAM [76] does not consider maintaining the magnitude consistency of the Softmax outputs changes drastically when input resolutions vary during testing. In contrast, our NSF explicitly maintains the magnitude consistency and alleviates the performance degradation. Furthermore, by adding BM-AM, the performance could be further boosted when extrapolating to a larger test length. These improvements confirm that these individual components are complementary and together they significantly promote the performance.
@@
Impact of Slope in Bias-Mode Attention Mask. In the Bias-Mode Attention Mask, the magnitude of the slope $\beta$ determines penalty rates in different heads. In the zero-shot case, we use a fixed slope. Table 5 reveals the impact of different slopes of BM-AM on the ISIC [12] dataset. We found that the best performance is achieved when $\beta=0.1$. Besides, our method is robust to different slope choices. In addition to the zero-shot setting, we also validate the BM-AM in fine-tuning scenarios on four datasets, shown in Table 1. The learnable parameters brought more performance growth compared with the fixed slope, while also bringing a slight increase in computation.
Computational Efficiency. In Table 6, we analyze the computational efficiency between the baselines and our BA-SAM. All the experiments are conducted on the same NVIDIA RTX 4090GPU to ensure fair comparisons. From the table, we observe that our BA-SAM is highly lightweight, incurring negligible computational overhead for the models. The reasons lie in two aspects: firstly, the NSF exhibits nearly identical computational complexity with the vanilla one. Besides, the BM-AM is seamlessly incorporated by adding a mask matrix to the query-key dot product before applying the softmax operation. Although there is a slight increase in memory usage, it remains negligible compared to the memory occupied by large models.
@@
Size 0.1 0.5 1 1.5 2 2.5
MAE 0.0375 0.0383 0.0381 0.0393 0.0388 0.0391
$\Delta$diff 40.4 40.8 40.4 40.9 41.2 40.7
Table 5: Impact of the slope $\beta$ in BM-AM (zero-shot setting).
@@ 6 Conclusion
In this paper, we address the important problem of varying image resolutions in SAM models by reformulating it as a problem of length extrapolation. To enhance the length extrapolation capability of SAM, we propose Scalable Bias-mode Attention Mask (BA-SAM). A new scaling factor (NSF) is introduced to maintain the consistent magnitude of attention. In addition, a bias-mode attention mask (BM-AM) is designed to prioritize neighboring information, mitigating the impact of untrained distant information. Extensive evaluation on diverse datasets reveals its ability to significantly alleviate performance degradation in the zero-shot setting and achieve state-of-the-art performance with minimal fine-tuning. Furthermore, we proposed a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets simultaneously.

### 19 Mar 2024 v3

@@ 4.2 Bias-Mode Attention Mask
Another challenge is that changes in token sequence length will lead to variations in positional encoding. It is important to ensure the insensitivity of the model when such positional encoding variations occur during the testing.
One possible way is absolute encoding without trainable parameters, such as Sinusoidal [62]. It requires the position encoding to have strong local-to-global inference capabilities. Nevertheless, this assumes that the given function has high-order smoothness (higher-order derivatives exist and are bounded). Commonly-used positional encodings are often combined with trigonometric functions These methods fail to satisfy the requirement of bounded high-order derivatives, making it less accurate to estimate the extrapolated results. Another potential approach is using local attention [47], which constrains the model’s field of view and remains insensitive to variations in token sequence length. However, local attention is typically implemented using a local window, necessitating modifications to the SAM structure, which requires re-training from scratch and is unacceptable for deployment in the downstream tasks.
To this end, we propose enabling the attention layer to focus more on the current token’s neighboring tokens. As such, even with an increase in the length of a token sequence, each token is scarcely affected by the untrained tokens from distant positions. In particular, we design a simple yet effective bias-mode mask. We introduce a bias after the query-key dot product.
As shown in Fig. 3, this mask exhibits a bias specified on the distance between the query-key pairs (i.e., $q\cdot k$). We expect that this proposed mask imposes penalties on attention scores between distant query-key pairs, and the penalty increases as the distance between a key $q$ and a query $k$ grows. To this end, we simply define the bias as the form of $b_{i,j}=\beta|i-j|$.
@@
We further discuss the setting of $\beta$ based on different cases. When conducting zero-shot generalization without fine-tuning, we set $\beta$ to a static, non-learned fixed value. The experimental section will discuss the specific value setting (Sec. 5). When fine-tuning is required, we make $\beta$ trainable. Since our Bias-Mode Attention Mask is lightweight relative to the model structure, it incurs negligible training cost overhead.
@@ 4.3 BA-SAM Model
As shown in Fig. 3, our BA-SAM is simple to implement, and can be seamlessly integrated into SAM [36] and its variants. Specifically, our design involves a new scaling factor (NSF) for the attention layer and a bias-mode attention mask (BM-AM). Our method does not involve any alterations to the model structure and is suited to both fine-tuning and non-fine-tuning cases. As for the fine-tuning case, it introduces negligible computational overhead as BM-AM incurs a very small amount of computation.
@@ 5 Experiments
Table 1: Performance comparisons in varying image resolutions. We employed the widely-used MAE (Mean Absolute Error) score. Lower MAE scores indicate better model performance. $\Delta$diff denotes the performance degradation due to resolution changes. Compared to the SAM [36] and MobileSAM [78] baselines, our proposed BA-SAM achieves smaller degradation when encountering token sequence length changes. The best MAE performance is highlighted in bold, and the smallest performance degradation is underlined.
@@ 5.1 Datasets and Implementations
Datasets. For a comprehensive evaluation of BA-SAM, we conduct extensive experiments on a wide range of segmentation tasks, i.e., salient object segmentation [18, 15, 20], complex object segmentation [19], skin lesion segmentation [96], camouflaged object detection [17], which correspond to four datasets: DUTS [64], DIS-TE4 [53], ISIC [12] and COD10K [16]. Besides, we verify it on the challenging COCO [45] instance segmentation benchmark. More details are referred to in the supplementary material.
Implementation Details. In zero-shot settings that do not require fine-tuning, we use the original SAM [36] backbone. For fine-tuning scenarios, we employ MobileSAM [78] as the backbone. MobileSAM [78] is a SAM variant with a structure similar to ViT-Tiny [73], and further details can be found in [78]. MobileSAM [78] uses a ViT-H-based SAM as a teacher network for distillation, ultimately achieving competitive accuracy compared to the original SAM but with significantly fewer parameters. For various object segmentation tasks, a random point is extracted from the ground truth as the prompt input during the fine-tuning phase. For instance segmentation, we use the ViT-B [14] backbone and the state-of-the-art detector Deformable-DETR  [94] trained on the COCO [45] dataset with Swin-L [47] backbone as box prompt generator. More details are provided in the supplementary material.
@@ 5.2 Results
Results of Various Object Segmentation Tasks: Tab. 1 demonstrates the effectiveness of our approach across four diverse segmentation datasets. $\Delta$diff denotes the value of the performance degradation due to resolution changes during the inference. The upper and lower parts of the table indicate the results without and with fine-tuning. The best MAE performance is highlighted in bold, and the smallest degradation is underlined. We have three observations: Firstly, our proposed BA-SAM consistently outperforms both SAM [36] and MobileSAM [78] baselines on all four datasets. This is mainly because these baselines do not consider the issue of varying image resolutions. In contrast, our presented scaling factor and bias-mode attention mask explicitly handle this issue and further alleviate the performance degradation. Secondly, when testing on higher resolutions than the training size, SAM [36] and MobileSAM [78] baselines show less desirable results than the original image size. In contrast, our BA-SAM incurs significantly less performance drop in different datasets. Thirdly, during the experiment, we observe negligible computational overhead, whether fine-tuning is applied or not, which supports the claim in the method section. See Sec. 5.3 for details.
Results of Instance Segmentation: In Tab. 3, we evaluate the performance of our method on the COCO [45] instance segmentation benchmark. For a fair comparison, all experiments are conducted in a zero-shot manner, with the same initialization parameters for the comparative methods and without the use of any additional training data. Our BA-SAM consistently outperforms SAM [36] and MobileSAM [78] baselines, demonstrating better zero-shot generalization capability on instance segmentation.
Comparisons with State-of-the-Art Methods: To further demonstrate the superiority of the effectiveness and generalizability of our method, we compare the state-of-the-art approaches in Tab. 2. From the table, we have two following observations: Firstly, all the state-of-the-art approaches [57, 79, 46, 95, 83, 53, 35, 16] show less-desirable performances in each dataset. Instead, our BA-SAM (specialized models) consistently outperforms these methods when fine-tuned on each downstream dataset. Secondly, almost all of these state-of-the-art techniques are specifically designed for one task and cannot be generalized well to other tasks. Due to the strong zero-shot generalization capability of SAM [36], our proposed BA-SAM can also be employed as a generalized model, which fine-tunes with all of these downstream datasets in a unified and shared model. Importantly, unlike [46, 57, 79, 12], we eliminate the need for employing additional techniques to further enhance the performance. As shown in Tab. 2, our generalized model also consistently promotes the performance of SAM on all datasets, demonstrating its remarkable generalizability.
@@ 5.3 Ablation Study and Analysis
Ablations Studies of Each Component. Tab. 4 summarizes the effect of each designed component on the settings with and without fine-tuning, respectively. The baseline means using the MobileSAM [78] as the base network that uses the vanilla scaling factor (VSF) in the attention layer [62]. New Scaling Factor and Bias-Mode Attention Mask are abbreviated as NSF and BM-AM, respectively. From the table, we observe that NSF could achieve superior performances than the baseline with VSF. This is because the vanilla attention in SAM [36] and MobileSAM [78] does not consider maintaining the magnitude consistency when Softmax outputs change drastically due to input resolutions vary during the testing. In contrast, our NSF explicitly maintains the magnitude consistency and alleviates the performance degradation. Furthermore, by adding BM-AM, the performance could be further boosted when extrapolating to a larger test length. These improvements confirm that these individual components are complementary and together they significantly promote the performance.
@@
Impact of Slope in Bias-Mode Attention Mask. In the Bias-Mode Attention Mask, the magnitude of the slope $\beta$ determines penalty rates in different heads. We found that the best performance is achieved when $\beta=0.1$. Besides, our method is robust to different slope choices. In the zero-shot case, we use a fixed slope $\beta=1$ by default. (More details could be referred to in the supplementary material)
Computational Efficiency. In Tab. 5, we analyze the computational efficiency between the baselines and our BA-SAM. All the experiments are conducted on the same NVIDIA RTX 4090GPU to ensure fair comparisons. From the table, we observe that our BA-SAM is highly lightweight, incurring negligible computational overhead for the models. The reasons lie in two aspects: firstly, the NSF exhibits nearly identical computational complexity with the vanilla one. Besides, the BM-AM is seamlessly incorporated by adding a mask matrix to the query-key dot product before applying the Softmax operation. Although there is a slight increase in memory usage, it remains negligible compared to the memory occupied by large models.
@@ 6 Conclusion
In this paper, we address the important problem of varying image resolutions in SAM models by reformulating it as a problem of length extrapolation. To enhance the length extrapolation capability of SAM, we propose the Scalable Bias-mode Attention Mask (BA-SAM) SAM. A new scaling factor (NSF) is introduced to maintain the consistent magnitude of attention. In addition, a bias-mode attention mask (BM-AM) is designed to prioritize neighboring information, mitigating the impact of untrained distant information. Extensive evaluation of diverse datasets reveals its ability to significantly alleviate performance degradation in the zero-shot setting and achieve state-of-the-art performance with minimal fine-tuning. Furthermore, we proposed a generalized model and benchmark, showcasing BA-SAM’s generalizability across all four datasets.
7 Acknowledgement
Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), Shanghai Science and Technology Commission (21511101200), National Natural Science Foundation of China (No. 72192821), YuCaiKe [2023] Project Number: 14105167-2023